OpenStudy (anonymous):

The ages of two persons differ by 16 years. If 6 years ago, the elder one be 3 times as old as the younger one, find their present ages

OpenStudy (anonymous):

there?

OpenStudy (anonymous):

yeah..

OpenStudy (anonymous):

have u tried something?

OpenStudy (anonymous):

Let the age of the younger person be x years. Then, age of the elder person = (x + 16) years. :. 3 (x - 6) = (x + 16 - 6) >> 3x -18 = x + 10 >> 2x = 28 Hence, their present ages are 14 years and 30 years.

OpenStudy (anonymous):

thats correct. :)